OpenStudy (anonymous):

How the heck do I write a set builder and interval out of 6-8(y+3)+5y>(2y-5)+13? once I worked it out I got y<-6m whats my next step?

OpenStudy (anonymous):

where did the m come from in your answer?

OpenStudy (anonymous):

thats a typo, I'm sorry.

OpenStudy (anonymous):

y<-6

OpenStudy (anonymous):

\[let s = \left\{ y: y<-6 \right\}\]

OpenStudy (anonymous):

in set builder, then the interval \[y \in \left( -6, -\infty \right)\]

OpenStudy (anonymous):

so how do I know where to put the infinity sign?

OpenStudy (anonymous):

oh wait i miss typed that, the -6 and - infinity shud be swapped. sorry

OpenStudy (anonymous):

the smaller value should always be on the left by convention since the interval starts at the left hand value and goes up to the right hand value

OpenStudy (anonymous):

so when do I use that squiggly line, and bracets instead of perenthisis?

OpenStudy (anonymous):

round brackets mean the endpoints arent in the interval and square brackets mean they are included

OpenStudy (anonymous):

so whats the interval to this specific one? \[\xi x-6\]

OpenStudy (anonymous):

im not sure i understand your notation, what does the lowercase delta stand for?

OpenStudy (anonymous):

Our professor was asking for some squiggly line for example: 8-6(x-3)>-4x+12 Solution: x<7 Set Builder Notation:\[\xi x-1+<7\xi\] Interval Notation: (-\[\infty , 5/2]\]

OpenStudy (anonymous):

I'm sorry the interval notation is actually 7 not 5/2, but do you see what I mean?

OpenStudy (anonymous):

maybe its just the name he is using but in any book i have ever seen set builder means { } and u state the conditions for membership of the set in the brackets

OpenStudy (anonymous):

yes, maybe it is {

OpenStudy (anonymous):

so to graph y<-6 it would be <_______O I I I I I I I I I I -6